# Towards verifications of Krylov complexity

### v1

@@ 5 Evaluation of $\mu_{2m}$ of exactly solvable discrete quantum mechanics
$=\sum_{n\in\mathcal{X}}\left\{\langle n|\eta^{2}|n\rangle\mathcal{A}_{m}(\mathcal{H})|n\rangle+\langle n|\eta\mathcal{L}\eta|n\rangle\mathcal{B}_{m}(\mathcal{H})|n\rangle+\langle n|\eta|n\rangle\mathcal{C}_{m}(\mathcal{H})|n\rangle\right\}/|\!|\eta|\!|^{2},$ (5.1)
all the coefficient terms $\mathcal{A}_{m}(\mathcal{H})|n\rangle$, $\mathcal{B}_{m}(\mathcal{H})|n\rangle$, $\mathcal{C}_{m}(\mathcal{H})|n\rangle$ are known
$\mathcal{H}|n\rangle=\mathcal{E}(n)|n\rangle\Rightarrow\ \alpha_{\pm}(\mathcal{H})|n\rangle=\alpha_{\pm}\bigl{(}\mathcal{E}(n)\bigr{)}|n\rangle,\quad R_{i}(\mathcal{H})|n\rangle=R_{i}\bigl{(}\mathcal{E}(n)\bigr{)}|n\rangle,\quad i=-1,0,1.$
@@
$\Longrightarrow\quad\langle n|\eta|n+1\rangle^{2}=A_{n}C_{n+1},\quad\langle n|\eta|n-1\rangle^{2}=A_{n-1}C_{n},\quad n\geq 1.$ (5.6)
@@
$|\!|\eta|\!|^{2}\mu_{m}$ $=\ \sum_{n\in\mathcal{X}}A_{n}C_{n+1}\Bigl{(}A_{m}\bigl{(}\mathcal{E}(n)\bigr{)}+\alpha_{+}\bigl{(}\mathcal{E}(n)\bigr{)}B_{m}\bigl{(}\mathcal{E}(n)\bigr{)}\Bigr{)}$
$\ \ +\sum_{n\in\mathcal{X}}A_{n-1}C_{n}\Bigl{(}A_{m}\bigl{(}\mathcal{E}(n)\bigr{)}+\alpha_{-}\bigl{(}\mathcal{E}(n)\bigr{)}B_{m}\bigl{(}\mathcal{E}(n)\bigr{)}\Bigr{)}.$
@@
$=\sum_{n=0}^{N-1}A_{n}C_{n+1}\left(\alpha_{+}\bigl{(}\mathcal{E}(n)\bigr{)}^{m}+\Bigl{(}-\alpha_{+}\bigl{(}\mathcal{E}(n)\bigr{)}\Bigr{)}^{m}\right)$
$=2\sum_{n=0}^{N-1}A_{n}C_{n+1}\left(\alpha_{+}\bigl{(}\mathcal{E}(n)\bigr{)}\right)^{m}\times\left\{\begin{array}[]{cc}1&m:\ \text{even}\\ 0&m:\ \text{odd}\end{array}\right.,$
@@
In the rest of this section, the necessary data for the evaluation of the moments $\mu_{2m}$ of the ten discrete quantum systems are provided. They are $B(x)$ and $D(x)$ for the definition of the Hamiltonian $\mathcal{H}$ and for specifying the parameter ranges, the sinusoidal coordinate $\eta(x)$, the energy eigenvalues $\mathcal{E}(n)$, $R_{0}(\mathcal{H})$ and $R_{1}(\mathcal{H})$ for the derivation of $\alpha_{\pm}\bigl{(}\mathcal{E}(n)\bigr{)}$ and the coefficients of the three term recurrence relation $A_{n}$ and $C_{n}$. It should be stressed that for these exactly solvable systems, the inside of the square root $R_{1}\bigl{(}\mathcal{E}(n)\bigr{)}^{2}+4R_{0}\bigl{(}\mathcal{E}(n)\bigr{)}$ (4.9) is always a complete square and $\alpha_{\pm}\bigl{(}\mathcal{E}(n)\bigr{)}$ are polynomials in $n$ or $q^{\pm n}$ of maximal degree two. For the full details of these exactly solvable discrete quantum systems, a paper by Odake and myself [14] should be consulted.
@@ 5.1 Krawtchouk and dual Hahn
$R_{0}(\mathcal{H})=(q^{-\tfrac{1}{2}}-q^{\tfrac{1}{2}})^{2}(H^{\prime})^{2},\quad R_{1}(\mathcal{H})=(q^{-\tfrac{1}{2}}-q^{\tfrac{1}{2}})^{2}H^{\prime},\quad\mathcal{H}^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{H}-1,\quad\mathcal{E}(n)=1-q^{n},$
@@ 5.4 $q$–Krawtchouk
×$\hskip 128.0374pt\times\frac{(1-q^{n+1})(1+pq^{n+1+N})}{(1+pq^{2n+1})(1+pq^{2n+2})}(q^{-n}+pq^{n+1})^{2m}$ (5.20)
@@ 6 Evaluation of $\mu_{m}$ of other exactly solvable quantum mechanical systems
In order to evaluate the moments $\{\mu_{m}\}$ for verifications of Krylov complexity of general quantum mechanical systems through orthonormalisation of a Krylov subspace, an appropriate definition of the inner product of operators is essential. The simplest trace one $(\mathcal{V},\mathcal{W})=\text{Tr}[\![\mathcal{V}^{\dagger}\mathcal{W}]\!]$ (2.8) is obviously ill-defined for most operators, which are unbounded. Here I tentatively adopt the one including the finite temperature ($T$) effects [5],
$(\mathcal{V},\mathcal{W})_{\beta}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{Z}\text{Tr}[\![e^{-\beta\mathcal{H}/2}\mathcal{V}^{\dagger}e^{-\beta\mathcal{H}/2}\mathcal{W}]\!]=\frac{1}{Z}\sum_{n=0}^{\infty}\langle n\left|e^{-\beta\mathcal{H}/2}\mathcal{V}^{\dagger}e^{-\beta\mathcal{H}/2}\mathcal{W}\right|n\rangle$ (6.1)
@@
However, as shown later in the examples §6.1–§6.3, the factor $Z$ cancels out in the calculation of the moments. The finite temperature inner product suppresses the contributions of higher energy eigenstates and make the trace summable. The norm of an operator is defined by
$(\mathcal{V},\mathcal{W})_{\beta}=(\mathcal{W},\mathcal{V})_{\beta}^{*}\rightarrow|\!|\mathcal{V}|\!|_{\beta}^{2}\stackrel{{\scriptstyle\text{def}}}{{=}}(\mathcal{V},\mathcal{V})_{\beta},$ (6.2)
With this inner product the flip property $(\mathcal{V},\mathcal{L}\mathcal{W})_{\beta}=(\mathcal{L}\mathcal{V},\mathcal{W})_{\beta}$ (2.10) holds but $\mathcal{V}^{\dagger}=\pm\mathcal{V}$ does not mean $(\mathcal{V},\mathcal{L}\mathcal{V})_{\beta}=0$. The Lanczos orthonormalisation reviewed in section 2 must be modified as follows [3]. At each step, the $\mathcal{O}_{k}$ component in $\mathcal{L}\mathcal{O}_{k}$ must be subtracted. One starts with
$\mathcal{O}_{0}\stackrel{{\scriptstyle\text{def}}}{{=}}\eta/|\!|\eta|\!|_{\beta},\quad(\mathcal{O}_{0},\mathcal{O}_{0})_{\beta}=1,$ (6.3)
$\mathcal{W}_{0}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{L}\mathcal{O}_{0}-a_{0}\mathcal{O}_{0},\quad a_{0}\stackrel{{\scriptstyle\text{def}}}{{=}}(\mathcal{O}_{0},\mathcal{L}\mathcal{O}_{0})_{\beta}\quad b_{1}\stackrel{{\scriptstyle\text{def}}}{{=}}|\!|\mathcal{W}_{0}|\!|_{\beta}\quad\mathcal{O}_{1}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{b_{1}}\mathcal{W}_{0}.$ (6.4)
Now $\mathcal{O}_{1}$ is orthogonal with the previous $\mathcal{O}_{0}$,
$(\mathcal{O}_{j},\mathcal{O}_{l})_{\beta}=\delta_{j\,l},\quad j,l=0,1.$ (6.5)
$\mathcal{W}_{1}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{L}\mathcal{O}_{1}-a_{1}\mathcal{O}_{1}-b_{1}\mathcal{O}_{0},\quad a_{1}\stackrel{{\scriptstyle\text{def}}}{{=}}(\mathcal{O}_{1},\mathcal{L}\mathcal{O}_{1})_{\beta}\quad b_{2}\stackrel{{\scriptstyle\text{def}}}{{=}}|\!|\mathcal{W}_{1}|\!|_{\beta},\quad\mathcal{O}_{2}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{b_{2}}\mathcal{W}_{1}.$ (6.6)
The orthogonality $(\mathcal{O}_{1},\mathcal{W}_{1})_{\beta}=0$ is obvious. By construction, the orthogonality $(\mathcal{O}_{0},\mathcal{W}_{1})_{\beta}=0$ holds. As
$(\mathcal{O}_{0},\mathcal{W}_{1})_{\beta}$ $=(\mathcal{O}_{0},\mathcal{L}\mathcal{O}_{1})_{\beta}-b_{1}(\mathcal{O}_{0},\mathcal{O}_{0})_{\beta}=(\mathcal{L}\mathcal{O}_{0},\mathcal{O}_{1})_{\beta}-b_{1}=\frac{1}{b_{1}}(\mathcal{W}_{0},\mathcal{W}_{0})_{\beta}-b_{1}=0,$
the orthonormality up to 2 is established
$(\mathcal{O}_{j},\mathcal{O}_{l})_{\beta}=\delta_{j\,l},\quad j,l=0,1,2.$ (6.7)
The flip property $(\mathcal{V},\mathcal{L}\mathcal{W})_{\beta}=(\mathcal{L}\mathcal{V},\mathcal{W})_{\beta}$ (2.10) plays an important role in the above and further calculations. The process goes on as
$\mathcal{W}_{k}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{L}\mathcal{O}_{k}-a_{k}\mathcal{O}_{k}-b_{k}\mathcal{O}_{k-1},\ a_{k}\stackrel{{\scriptstyle\text{def}}}{{=}}(\mathcal{O}_{k},\mathcal{L}\mathcal{O}_{k})_{\beta},\ b_{k+1}\stackrel{{\scriptstyle\text{def}}}{{=}}|\!|\mathcal{W}_{k}|\!|_{\beta},\ \mathcal{O}_{k+1}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{b_{k+1}}\mathcal{W}_{k}.$ (6.8)
It is easy to prove that $\mathcal{O}_{k+1}$ is orthogonal to all the previous ones $(\mathcal{O}_{j},\mathcal{O}_{k+1})_{\beta}=0$, $j=0,1,\ldots,k$, by assuming the previous ones are orthonormal
$(\mathcal{O}_{j},\mathcal{O}_{l})_{\beta}=\delta_{j\,l},\quad j,l=0,1,\ldots,k.$ (6.9)
It is obvious by construction $(\mathcal{O}_{k},\mathcal{W}_{k})_{\beta}=0=(\mathcal{O}_{k},\mathcal{O}_{k+1})_{\beta}$. Likewise,
$(\mathcal{O}_{k-1},\mathcal{W}_{k})_{\beta}$ $=\bigl{(}\mathcal{O}_{k-1},\mathcal{L}\mathcal{O}_{k}-b_{k}\mathcal{O}_{k-1}\bigr{)}_{\beta}=\bigl{(}\mathcal{O}_{k-1},\mathcal{L}\mathcal{O}_{k}\bigr{)}_{\beta}-b_{k}=\frac{1}{b_{k}}\bigl{(}\mathcal{O}_{k-1},\mathcal{L}\mathcal{W}_{k-1}\bigr{)}_{\beta}-b_{k}$
$=\frac{1}{b_{k}}\bigl{(}\mathcal{L}\mathcal{O}_{k-1},\mathcal{W}_{k-1}\bigr{)}_{\beta}-b_{k}=\frac{1}{b_{k}}\bigl{(}\mathcal{W}_{k-1}+b_{k-1}\mathcal{O}_{k-2},\mathcal{W}_{k-1}\bigr{)}_{\beta}-b_{k}=b_{k}-b_{k}=0,$
as $(\mathcal{O}_{k-2},\mathcal{W}_{k-1})_{\beta}=b_{k}(\mathcal{O}_{k-2},\mathcal{O}_{k})_{\beta}=0$ by assumption. For $j\leq k-2$, the same logic goes
$(\mathcal{O}_{j},\mathcal{W}_{k})_{\beta}$ $=\bigl{(}\mathcal{O}_{j},\mathcal{L}\mathcal{O}_{k}-b_{k}\mathcal{O}_{k-1}\bigr{)}_{\beta}=(\mathcal{O}_{j},\mathcal{L}\mathcal{O}_{k})_{\beta}=\frac{1}{b_{k}}(\mathcal{L}\mathcal{O}_{j},\mathcal{W}_{k-1})_{\beta}$
$=\frac{1}{b_{k}}(\mathcal{W}_{j}+b_{j}\mathcal{O}_{j-1},\mathcal{W}_{k-1})_{\beta}=0,$
and the induction is complete. $\square$
@@
$\mathcal{L}^{m}\eta=$ $\,\eta\,\mathcal{A}_{m}(\mathcal{H})+\mathcal{L}\eta\,\mathcal{B}_{m}(\mathcal{H})+\mathcal{C}_{m}(\mathcal{H}),\qquad m\in\mathbb{N}_{\geq 0},$ (4.4)
@@
Due to the property $(\mathcal{V},\mathcal{L}\mathcal{V})_{\beta}\neq 0$ of the finite temperature trace, odd moments as well as even moments
$\mu_{m}=(\mathcal{O}_{0},\mathcal{L}^{m}\mathcal{O}_{0})_{\beta}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{Z|\!|\eta|\!|_{\beta}^{2}}\text{Tr}[\![e^{-\beta\mathcal{H}/2}\eta e^{-\beta\mathcal{H}/2}\mathcal{L}^{m}\eta]\!]$ (6.10)
need be evaluated. Since the denominator $Z|\!|\eta|\!|_{\beta}^{2}$ is independent of $m$, the numerator only
@@
The orthonormal eigenvectors of the Hamiltonian $\mathcal{H}$ $|n\rangle$ have the same structure as (3.3) and (3.4). But the forms of the sinusoidal coordinates $\eta(x)$ and the energy spectrum $\mathcal{E}(n)$ are not restricted to those listed in (3.5) and (3.6). As the types of the eigenpolynomials $\{P_{n}(\eta)\}$ are more varied than those of the discrete quantum mechanics, the uniform normalisatin condition $P_{n}(0)=1$ (3.7) no longer applies and the three term recurrence relation has a more general form
@@
$\langle n|\eta|n\rangle\langle n|\mathcal{L}^{m}\eta|n\rangle=\langle n|\eta|n\rangle\left\{\langle n|\eta|n\rangle\mathcal{A}_{m}\bigl{(}\mathcal{E}(n)\bigr{)}+\langle n|\mathcal{L}\eta|n\rangle\mathcal{B}_{m}\bigl{(}\mathcal{E}(n)\bigr{)}+\mathcal{C}_{m}\bigl{(}\mathcal{E}(n)\bigr{)}\right\},$
@@
$=\sum_{n=0}^{\infty}\sum_{i=n,n\pm 1}e^{-\beta\mathcal{E}(n)/2}e^{-\beta\mathcal{E}(l)/2}\langle n|\eta|l\rangle\langle l|\eta|n\rangle$
@@
$\mu_{2m}=\frac{2}{Z|\!|\eta|\!|_{\beta}^{2}}\sum_{n=0}^{N-1}e^{-\beta(\mathcal{E}(n)+\mathcal{E}(n+1))/2}A_{n}C_{n+1}\left(\alpha_{+}\bigl{(}\mathcal{E}(n)\bigr{)}\right)^{2m},\quad\mu_{2m-1}=0,\quad m\in\mathbb{N}.$ (6.21)
The vanishing of the odd moments is the consequence of the exact solvability of the Heisenberg equation of motion (4.4), (4.14). It applies to two non-compact discrete quantum systems with the eigenpolynomials, the Meixner (M) and Charlier (C), as well as the exactly solvable systems related to the Hermite (H), Laguerre (L), Gegenbauer (G), and Jacobi (J) polynomials.
Remark 6.2
The vanishing of the odd moments $\mu_{2m-1}=0$ simply means that all $a_{k}=(\mathcal{O}_{k},\mathcal{L}\mathcal{O}_{k})_{\beta}$ in (6.8) are also vanishing. Therefore, the relationship between the Lanczos coefficients $\{b_{n}\}$ and the moments $\{\mu_{2m}\}$ mentioned in Remark 2.4 is also valid with the trace including the finite temperature effects (6.1).
In the rest of this section, the necessary data for the evaluation of the moments $\mu_{2m}$ of the above six exactly solvable quantum mechanical systems are provided. For the ordinary one-dimensional quantum systems, they are the potential $U(x)$ in the Hamiltonian $\mathcal{H}=p^{2}+U(x)$ in which $p$ is the canonical momentum operator conjugate to $x$, $[p,x]=-i$, the sinusoidal coordinate $\eta(x)$, the energy eigenvalues $\mathcal{E}(n)$, $R_{0}(\mathcal{H})$ and $R_{1}(\mathcal{H})$ for the derivation of $\alpha_{\pm}\bigl{(}\mathcal{E}(n)\bigr{)}$ and the coefficients of the three term recurrence relation $A_{n}$ $B_{n}$ and $C_{n}$.

### v2

@@ 5 Evaluation of $\mu_{2m}$ of exactly solvable discrete quantum mechanics
$=\sum_{n\in\mathcal{X}}\left\{\langle n|\eta^{2}|n\rangle\langle n|\mathcal{A}_{m}(\mathcal{H})|n\rangle+\langle n|\eta\mathcal{L}\eta|n\rangle\langle n|\mathcal{B}_{m}(\mathcal{H})|n\rangle+\langle n|\eta|n\rangle\langle n|\mathcal{C}_{m}(\mathcal{H})|n\rangle\right\}/|\!|\eta|\!|^{2},$ (5.1)
all the coefficient terms $\langle n|\mathcal{A}_{m}(\mathcal{H})|n\rangle$, $\langle n|\mathcal{B}_{m}(\mathcal{H})|n\rangle$, $\langle n|\mathcal{C}_{m}(\mathcal{H})|n\rangle$ are known
$\mathcal{H}|n\rangle=\mathcal{E}(n)|n\rangle\Rightarrow\ \langle n|\alpha_{\pm}(\mathcal{H})|n\rangle=\alpha_{\pm}\bigl{(}\mathcal{E}(n)\bigr{)},\quad\langle n|R_{i}(\mathcal{H})|n\rangle=R_{i}\bigl{(}\mathcal{E}(n)\bigr{)},\quad i=-1,0,1.$
@@
$\Longrightarrow\quad\langle n|\eta|n+1\rangle^{2}=A_{n}C_{n+1},\quad n\leq N-1,\quad\langle n|\eta|n-1\rangle^{2}=A_{n-1}C_{n},\quad n\geq 1.$ (5.6)
@@
$|\!|\eta|\!|^{2}\mu_{m}$ $=\ \sum_{n=0}^{N-1}A_{n}C_{n+1}\Bigl{(}A_{m}\bigl{(}\mathcal{E}(n)\bigr{)}+\alpha_{+}\bigl{(}\mathcal{E}(n)\bigr{)}B_{m}\bigl{(}\mathcal{E}(n)\bigr{)}\Bigr{)}$
$\ \ +\sum_{n=1}^{N}A_{n-1}C_{n}\Bigl{(}A_{m}\bigl{(}\mathcal{E}(n)\bigr{)}+\alpha_{-}\bigl{(}\mathcal{E}(n)\bigr{)}B_{m}\bigl{(}\mathcal{E}(n)\bigr{)}\Bigr{)}.$
@@
$=\sum_{n=0}^{N-1}A_{n}C_{n+1}\Bigl{(}\alpha_{+}\bigl{(}\mathcal{E}(n)\bigr{)}^{m}+\bigl{(}-\alpha_{+}\bigl{(}\mathcal{E}(n)\bigr{)}\bigr{)}^{m}\Bigr{)}$
$=2\sum_{n=0}^{N-1}A_{n}C_{n+1}\Bigl{(}\alpha_{+}\bigl{(}\mathcal{E}(n)\bigr{)}\Bigr{)}^{m}\times\left\{\begin{array}[]{cc}1&m:\ \text{even}\\ 0&m:\ \text{odd}\end{array}\right.,$
@@
In the rest of this section, the necessary data for the evaluation of the moments $\mu_{2m}$ of the ten discrete quantum systems are provided. They are $B(x)$ and $D(x)$ for the definition of the Hamiltonian $\mathcal{H}$ and for specifying the parameter ranges, the sinusoidal coordinates $\eta(x)$, the energy eigenvalues $\mathcal{E}(n)$, $R_{0}(\mathcal{H})$ and $R_{1}(\mathcal{H})$ for the derivation of $\alpha_{\pm}\bigl{(}\mathcal{E}(n)\bigr{)}$ and the coefficients of the three term recurrence relation $A_{n}$ and $C_{n}$. It should be stressed that for these exactly solvable systems, the inside of the square root $R_{1}\bigl{(}\mathcal{E}(n)\bigr{)}^{2}+4R_{0}\bigl{(}\mathcal{E}(n)\bigr{)}$ (4.9) is always a complete square and $\alpha_{\pm}\bigl{(}\mathcal{E}(n)\bigr{)}$ are polynomials in $n$ or $q^{\pm n}$ of maximal degree two. For the full details of these exactly solvable discrete quantum systems, a paper by Odake and myself [16] should be consulted.
@@ 5.1 Krawtchouk and dual Hahn
$R_{0}(\mathcal{H})=(q^{-\tfrac{1}{2}}-q^{\tfrac{1}{2}})^{2}(\mathcal{H}^{\prime})^{2},\quad R_{1}(\mathcal{H})=(q^{-\tfrac{1}{2}}-q^{\tfrac{1}{2}})^{2}\mathcal{H}^{\prime},\quad\mathcal{H}^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{H}-1,\quad\mathcal{E}(n)=1-q^{n},$
@@ 5.4 $q$–Krawtchouk
×$\hskip 128.0374pt\times\frac{(1-q^{n+1})(1+pq^{n+1+N})}{(1+pq^{2n+1})(1+pq^{2n+2})}(q^{-n}+pq^{n+1})^{2m},$ (5.20)
@@ 6 Evaluation of $\mu_{m}$ of other exactly solvable quantum mechanical systems
In order to evaluate the moments $\{\mu_{m}\}$ for verifications of Krylov complexity of general quantum mechanical systems through orthonormalisation of a Krylov subspace, an appropriate definition of the inner product of operators is essential. The simplest trace one $(\mathcal{V},\mathcal{W})=\text{Tr}[\![\mathcal{V}^{\dagger}\mathcal{W}]\!]$ (2.8) is obviously ill-defined for most operators, which are unbounded. A simple prescription to deal with unbounded operators is to introduce the finite temperature (T) effects through the Boltzmann factor $e^{-\beta\mathcal{H}}$, $\beta=1/T$. The finite temperature inner product suppresses the contributions of higher energy eigenstates and make the trace summable. The general forms of the inner product including the Boltzmann factor were discussed in detail in section VIII of [1] and [5]. Here I choose, following [1, 5], the Wightman inner product
$(\mathcal{V},\mathcal{W})_{\beta}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{Z}\text{Tr}[\![e^{-\beta\mathcal{H}/2}\mathcal{V}^{\dagger}e^{-\beta\mathcal{H}/2}\mathcal{W}]\!]=\frac{1}{Z}\sum_{n=0}^{\infty}\langle n\left|e^{-\beta\mathcal{H}/2}\mathcal{V}^{\dagger}e^{-\beta\mathcal{H}/2}\mathcal{W}\right|n\rangle,$ (6.1)
@@
However, as shown later in the examples §6.1–§6.3, the factor $Z$ cancels out in the calculation of the moments. The norm of an operator is defined by
$(\mathcal{V},\mathcal{W})_{\beta}=(\mathcal{W}^{\dagger},\mathcal{V}^{\dagger})_{\beta}=(\mathcal{W},\mathcal{V})_{\beta}^{*}\rightarrow|\!|\mathcal{V}|\!|_{\beta}^{2}\stackrel{{\scriptstyle\text{def}}}{{=}}(\mathcal{V},\mathcal{V})_{\beta},$ (6.2)
With this symmetric inner product the flip property $(\mathcal{V},\mathcal{L}\mathcal{W})_{\beta}=(\mathcal{L}\mathcal{V},\mathcal{W})_{\beta}$ (2.10) holds and $\mathcal{V}^{\dagger}=\pm\mathcal{V}$ means
$(\mathcal{V},\mathcal{L}\mathcal{V})_{\beta}=\bigl{(}(\mathcal{L}\mathcal{V})^{\dagger},\mathcal{V}^{\dagger})_{\beta}=(-\mathcal{L}\mathcal{V},\mathcal{V})_{\beta}=-(\mathcal{V},\mathcal{L}\mathcal{V})_{\beta}=0.$ (6.3)
Thus the Lanczos orthonormalisation reviewed in section 2 works equally for the Wightman inner product (6.1).
It should be stressed that the situation is different for non-symmetric inner products, for example [3],
$(\mathcal{V},\mathcal{W})_{\rho}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{Z}\text{Tr}[\![\mathcal{V}^{\dagger}\rho_{1}\mathcal{W}\rho_{2}]\!]=(\mathcal{W},\mathcal{V})_{\rho}^{*}\neq(\mathcal{W}^{\dagger},\mathcal{V}^{\dagger})_{\rho},$
in which $\rho_{1},\rho_{2}$ are positive functions of the Boltzmann factor. For $\rho_{1}\neq\rho_{2}$, the flip property $(\mathcal{V},\mathcal{L}\mathcal{W})_{\beta}=(\mathcal{L}\mathcal{V},\mathcal{W})_{\beta}$ (2.10) holds but $\mathcal{V}^{\dagger}=\pm\mathcal{V}$ does not mean $(\mathcal{V},\mathcal{L}\mathcal{V})_{\beta}=0$. Therefore the Lanczos orthonormalisation must be modified, as shown in [3].
@@
$\mathcal{L}^{m}\eta=$ $\,\eta\,\mathcal{A}_{m}(\mathcal{H})+\mathcal{L}\eta\,\mathcal{B}_{m}(\mathcal{H})+\mathcal{C}_{m}(\mathcal{H}),\qquad m\in\mathbb{Z}_{\geq 0},$ (4.4)
@@
Let us evaluate the moments
$\mu_{m}=(\mathcal{O}_{0},\mathcal{L}^{m}\mathcal{O}_{0})_{\beta}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{Z|\!|\eta|\!|_{\beta}^{2}}\text{Tr}[\![e^{-\beta\mathcal{H}/2}\eta e^{-\beta\mathcal{H}/2}\mathcal{L}^{m}\eta]\!],\qquad m\in\mathbb{Z}_{\geq 0}.$ (6.4)
Since the denominator $Z|\!|\eta|\!|_{\beta}^{2}$ is independent of $m$, the numerator only
@@
The orthonormal eigenvectors of the Hamiltonian $\mathcal{H}$, $\{|n\rangle\}$ have the same structure as (3.3) and (3.4). But the forms of the sinusoidal coordinates $\eta(x)$ and the energy spectrum $\mathcal{E}(n)$ are not restricted to those listed in (3.5) and (3.6). As the types of the eigenpolynomials $\{P_{n}(\eta)\}$ are more varied than those of the discrete quantum mechanics, the uniform normalisation condition $P_{n}(0)=1$ (3.7) no longer applies and the three term recurrence relation has a more general form
@@
$\langle n|\eta|n\rangle\langle n|\mathcal{L}^{m}\eta|n\rangle=\langle n|\eta|n\rangle\left\{\langle n|\eta|n\rangle\mathcal{A}_{m}\bigl{(}\mathcal{E}(n)\bigr{)}+\langle n|\mathcal{L}\eta|n\rangle\mathcal{B}_{m}\bigl{(}\mathcal{E}(n)\bigr{)}+\mathcal{C}_{m}\bigl{(}\mathcal{E}(n)\bigr{)}\right\}=0,$
@@
$=\sum_{n=0}^{\infty}\sum_{l=n,n\pm 1}e^{-\beta\mathcal{E}(n)/2}e^{-\beta\mathcal{E}(l)/2}\langle n|\eta|l\rangle\langle l|\eta|n\rangle$
@@
$\hskip 142.26378pt\left.+e^{-\beta\bigl{(}\mathcal{E}(n)+\mathcal{E}(n-1)\bigr{)}/2}\langle n|\eta|n-1\rangle\langle n-1|\eta|n\rangle\right\}$
@@
$\mu_{2m}=\frac{2}{Z|\!|\eta|\!|_{\beta}^{2}}\sum_{n=0}^{\infty}e^{-\beta(\mathcal{E}(n)+\mathcal{E}(n+1))/2}A_{n}C_{n+1}\left(\alpha_{+}\bigl{(}\mathcal{E}(n)\bigr{)}\right)^{2m},\quad\mu_{2m-1}=0,\quad m\in\mathbb{N}.$ (6.15)
It applies to two non-compact discrete quantum systems with the eigenpolynomials, the Meixner (M) and Charlier (C), as well as the exactly solvable systems related to the Hermite (H), Laguerre (L), Gegenbauer (G), and Jacobi (J) polynomials.
In the rest of this section, the necessary data for the evaluation of the moments $\mu_{2m}$ of the above six exactly solvable quantum mechanical systems are provided. For the ordinary one-dimensional quantum systems, they are the potential $U(x)$ in the Hamiltonian $\mathcal{H}=p^{2}+U(x)$ in which $p$ is the canonical momentum operator conjugate to $x$, $[p,x]=-i$, the sinusoidal coordinates $\eta(x)$, the energy eigenvalues $\mathcal{E}(n)$, $R_{0}(\mathcal{H})$ and $R_{1}(\mathcal{H})$ for the derivation of $\alpha_{\pm}\bigl{(}\mathcal{E}(n)\bigr{)}$ and the coefficients of the three term recurrence relation $A_{n}$ $B_{n}$ and $C_{n}$.